OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2). y = −x − 6 y = −x − 14 y = 3x − 6 y = 3x − 14

OpenStudy (anonymous):

y = 2x + 3

OpenStudy (anonymous):

thats not in the choices

OpenStudy (anonymous):

srry.... parallel to the line 3x - y = -3 and passes through the point (4, -2). Use 3x - y = b Substitute (4, -2) into the equation to find b, 3(4) + 2 = b b = 14 Ans: 3x - y = 14 or y = 3x - 14. Select D.